# Explicit Abstention Knobs for Predictable Reliability in Video Question Answering

### v1

@@ 1.2 This Work
    Confidence-based abstention provides mechanistic control in-distribution. Sweeping threshold $\varepsilon$ produces smooth, monotone risk-coverage curves with a clear operating regime (63.7% coverage, 9.4% error vs. 99% coverage, 23.6% error).
@@ 3.1 Dataset
Each question has five multiple-choice options labeled A through E. We evaluate on 300 items from the validation split, stratified to include 100 questions of each type. This item list is frozen in item_ids.json to ensure exact reproducibility across all experiments, including the Evidence Degradation conditions.
@@ 3.2 Evidence Packet Construction
Parameter Baseline Shift A
@@ 3.5 Evaluation Metrics
We also measure calibration using Expected Calibration Error (ECE) (Guo et al., 2017), computed over the accepted predictions at each threshold. ECE quantifies the alignment between predicted confidence and empirical accuracy by binning predictions and measuring the weighted average of absolute differences between mean confidence and accuracy in each bin. Well-calibrated predictions should have ECE near zero.
@@ 3.6 Logprob-Derived Confidence
Or equivalently in log space (Figure 2):
$\log p_{i}=\ell_{i}-\log\sum_{k\in\{A,B,C,D,E\}}e^{\ell_{k}}$ (3)
Figure 2: Logprob-to-probability computation. Raw log probabilities $\ell_{i}$ from the model’s token distribution are normalized via log-softmax over the five answer options. This renormalization is necessary because the model’s full vocabulary distribution includes tokens beyond A–E; we restrict to the answer space and compute a proper probability distribution over only the valid options. The resulting $p_{i}$ values sum to 1 and represent the model’s relative preference among answer choices.
The key insight is that raw logprobs $\ell_{i}$ from the Vertex AI response reflect the model’s scoring over its entire vocabulary, not just the answer options. By extracting only the A–E logprobs and renormalizing, we obtain a probability distribution over the answer space that represents how the model’s token-level preference is distributed among the five choices. This is the decoder’s “voting distribution” over answers, independent of the self-reported confidence scalar.
@@ 4.3 Evidence Degradation
Figure 3: Visual comparison of evidence packets for the same video. Top row shows 6 frames sampled from the original 18-frame evidence packet. Bottom row shows all 6 frames from the degraded condition (Shift A). The degraded condition provides much sparser temporal coverage of the video’s 76.5-second duration.
@@ 4.4 Logprob Confidence
We query the same 300 items across four evidence conditions: original (18 frames), Shift A (6 frames), early-only (first 50% of video), and late-only (second 50% of video). For each query, we extract the token-level probability distribution over answer options and compute three confidence metrics: maximum probability ($p_{\max}$), margin, and normalized entropy.
@@ 5.1.2 Calibration Analysis
Figure 4: Baseline performance. (a) Risk-Coverage curve showing smooth tradeoff with knee at 60–70% coverage; starred point marks $\varepsilon=0.71$ (9.4% risk, 63.7% coverage). (b) Reliability diagram at $\varepsilon=0.71$ showing good calibration in the high-confidence bin. (c) ECE decreases as threshold tightens, dropping from 0.067 to below 0.02.
@@ 5.2 Evidence Degradation
Figure 5: Concrete example of overconfidence under evidence degradation. The question requires observing sustained behavior across time. With 18 frames (6 shown here), the model correctly identifies the behavior (B) with confidence 1.00. With only 6 frames from the degraded condition, the model misses critical moments, answers incorrectly (A), yet reports confidence 0.70. The model fails to recognize that sparse sampling provides insufficient evidence.
@@ 5.2.1 The Overconfidence Problem
Figure 6: Risk-Coverage comparison under evidence degradation. The Shift A curve (6 frames, red) is shifted leftward and slightly upward from the original (18 frames, blue). At any fixed epsilon, the 6-frame regime has lower coverage but similar conditional risk. Compression shift (green) has minimal impact.
@@ 5.3 Logprob Confidence
Shift A (6 frames) 81.3% 0.876 0.974 0.771 0.282
Early-only (6 frames) 77.3% 0.861 0.961 0.744 0.319
Late-only (6 frames) 78.7% 0.872 0.979 0.772 0.302
@@ 5.3.1 Comparison with Self-Reported Confidence
Table 5 directly compares self-reported confidence (from JSON output) with logprob-derived $p_{\max}$. The comparison reveals that logprob confidence is systematically higher and less responsive to evidence degradation. Figure 7 visualizes the risk-coverage curves for both confidence methods under evidence degradation.
Figure 7: Risk-coverage curves comparing self-reported confidence (left) and logprob-derived $p_{\max}$ (right) under evidence degradation. Both methods exhibit leftward shift when frames are reduced from 18 to 6. However, the gap between conditions is smaller for logprob confidence, indicating that the model’s token-level decision distribution is even less sensitive to evidence quality than self-reported confidence.
@@
Mean 0.828 0.786 0.871 0.876
Median 0.900 0.900 0.970 0.974
$\Delta$ (degradation) -5.1% ++0.6%
Table 5: Self-reported vs. logprob-derived confidence under evidence degradation. Self-reported confidence drops 5.1% (0.828 $\to$ 0.786) when frames are reduced. Logprob $p_{\max}$ actually increases slightly (0.871 $\to$ 0.876). The model’s token-level decision distribution is even less sensitive to evidence quality than its self-reported confidence.
@@ 5.3.2 Interpretation
  • Self-reported confidence drops 5.1% under degradation (0.828 $\to$ 0.786)
  • Logprob $p_{\max}$ increases 0.6% (0.871 $\to$ 0.876)
@@
The temporal ablation conditions (early-only and late-only) reinforce this conclusion. For questions whose answers depend on events in the missing temporal segment, these conditions provide semantically insufficient evidence by construction. Yet $p_{\max}$ remains above 0.86 and median $p_{\max}$ exceeds 0.96 in both cases. Because temporal ablation guarantees removal of one half of the clip, stability of $p_{\max}$ under early-only and late-only conditions suggests insensitivity is not explained solely by redundant sampling—the model genuinely does not recognize when critical temporal context is absent.
@@ 6.1 Summary of Findings
Third, the logprob analysis confirms that overconfidence is fundamental to the model’s representations, not an artifact of the self-reporting interface. Logprob-derived $p_{\max}$ actually increases slightly under evidence degradation (0.871 $\to$ 0.876), while self-reported confidence at least decreases modestly (0.828 $\to$ 0.786). The model’s token-level decision distribution over answer options does not become more diffuse when evidence quality degrades.
@@ 6.3.1 What These Experiments Support
These three facts imply a warrant-based formulation in which reported confidence $p$ should satisfy $p\leq\zeta(e)+\epsilon$ for some evidence-derived bound $\zeta(e)$. The Shift A failure is precisely the violation this contract would rule out. Confidence remains high when the evidence channel is weaker.
@@ 6.3.6 Observability Sensitivity Diagnostic
$\Pr(\text{conf}\geq 0.9)$ 63.0% (189/300) 53.3% (160/300)
$\Pr(\text{wrong}\mid\text{conf}\geq 0.9)$ 9.0% 8.7%
Mean confidence 0.828 0.786
@@
Confidence does not contract commensurate with evidence loss. Despite reducing frames from 18 to 6, median confidence remains identical at 0.900. The high-confidence rate drops modestly (63.0% to 53.3%), but the error rate among high-confidence predictions is nearly identical (9.0% vs 8.7%). Confidence is not evidence-conditioned. The model maintains high confidence despite reduced observability, and the modest selectivity increase (coverage drops 16% at fixed $\varepsilon$) is not proportional to the 67% evidence reduction.
@@ Appendix B Per-Category Degradation
We analyze degradation patterns across question categories (Causal: CW+CH, Temporal: TN+TC+TP, Descriptive: DO+DL+DC) at three operating points to prevent cherry-picking. Note that the overall coverage reported in Table 3 (e.g., 53.7% at $\varepsilon=0.71$ under Shift A) is the average across these three 100-item strata; the per-category breakdown below explains the aggregate behavior.
@@ Appendix C Temporal Ablation
For each video, we generate two 6-frame evidence packets: early-only (frames sampled uniformly from the first 50% of the clip) and late-only (frames sampled uniformly from the second 50%). This design targets temporal and causal questions, which often require observing sequences of events spanning the full video. By restricting frames to one half, we remove evidence about events occurring in the other half—a more controlled intervention on semantic content than uniform subsampling, which may still capture key moments regardless of density.
@@
Shift A (6 uniform) 97.3% 27.4% 72.6% 0.786
Early-only (0–50%) 97.3% 25.7% 74.3% 0.792
Late-only (50–100%) 97.7% 25.3% 74.7% 0.800
@@
Shift A (6 uniform) 53.7% 9.3% 90.7% 0.926
Early-only (0–50%) 52.3% 8.9% 91.1% 0.923
Late-only (50–100%) 51.7% 10.3% 89.7% 0.929
@@
Table 10: Temporal Ablation results. At $\varepsilon=0$, early-only and late-only conditions perform similarly to uniform 6-frame sampling (25–26% risk vs 27%). At $\varepsilon=0.71$, all 6-frame conditions converge to similar coverage (51–54%) with mean confidence remaining above 0.92. The model does not differentiate between early and late evidence despite the procedural removal of temporal context.
@@
Temporal Ablation conditions behave nearly identically to uniform 6-frame subsampling. At $\varepsilon=0$, early-only (25.7% risk) and late-only (25.3% risk) perform comparably to Shift A (27.4% risk). At $\varepsilon=0.71$, all three 6-frame conditions converge to similar coverage (51–54%) and mean confidence ($\sim$0.92). The model does not “notice” whether it is seeing the first or second half of the video. Confidence remains high regardless of which temporal segment is provided.
@@
This finding strengthens our claim that confidence does not track evidence completeness. Unlike uniform subsampling, which may preserve key moments by chance, temporal ablation guarantees removal of one half of the clip’s temporal context. For questions whose answers depend on late events (e.g., “what does X do after Y?”), early-only evidence is semantically insufficient. Yet the model maintains similarly high confidence in both conditions. This suggests overconfidence is not merely an artifact of temporal redundancy in the dataset, but reflects a fundamental insensitivity to evidence availability.
Appendix D Prompt Template

### v2

@@ 1.2 This Work
    Confidence-based abstention provides mechanistic control in-distribution. Sweeping threshold $\varepsilon$ produces smooth, monotone risk-coverage curves with a clear operating regime (63.7% coverage, 9.4% error vs. 98.7% coverage, 23.6% error).
@@ 3.1 Dataset
Each question has five multiple-choice options labeled A through E. We evaluate on 300 items from the validation split, stratified to include 100 questions of each type. This item list is frozen in item_ids.json to ensure exact reproducibility across all experiments, including the Evidence Degradation conditions. For any cross-method comparison (self-reported vs. logprob-derived confidence), we match predictions by question_id under the same evidence condition and report statistics on the intersection where both methods produce valid outputs ($n=295$ at 18 frames, $n=292$ at 6 frames).
@@ 3.2 Evidence Packet Construction
Parameter Baseline Sparse (6f)
@@ 3.5 Evaluation Metrics
We also measure calibration using Expected Calibration Error (ECE) (Guo et al., 2017), computed over the accepted predictions at each threshold:
$\text{ECE}=\sum_{b=1}^{B}\frac{|B_{b}|}{n}\left|\text{acc}(B_{b})-\text{conf}(B_{b})\right|$ (2)
where predictions are partitioned into $B$ bins by confidence, $B_{b}$ is the set of predictions in bin $b$, $\text{acc}(B_{b})$ is the fraction correct, and $\text{conf}(B_{b})$ is the mean confidence. Well-calibrated predictions should have ECE near zero.
@@ 3.6 Logprob-Derived Confidence
This renormalization is necessary because the model’s full vocabulary distribution includes tokens beyond A–E. By extracting only the A–E logprobs and renormalizing, we obtain a probability distribution over the answer space that represents the model’s token-level preference among the five choices—the decoder’s “voting distribution” over answers, independent of the self-reported confidence scalar.
@@
Importantly, self-reported confidence and logprob-derived confidence are obtained via separate model calls with different output interfaces (JSON vs. single-letter response with logprobs enabled). This can change instruction-following behavior and therefore absolute accuracy, so we avoid comparing cross-interface accuracy levels. Our comparisons focus on (i) how each confidence signal changes under evidence degradation (18 $\to$ 6 frames) within the same interface, and (ii) whether the qualitative failure mode—confidence not contracting under evidence loss—persists across interfaces. When we juxtapose the two confidence signals in a single figure or table, we compute statistics on matched question instances (same question_id, same evidence condition), restricted to the intersection where both calls produce valid outputs.
@@ 4.3 Evidence Degradation
Figure 2: Visual comparison of evidence packets for the same video. Top row shows 6 frames sampled from the original 18-frame evidence packet. Bottom row shows all 6 frames from the degraded condition (Sparse). The degraded condition provides much sparser temporal coverage of the video’s 76.5-second duration.
@@ 4.4 Logprob Confidence
We query the same 300 items across four evidence conditions: original (18 frames), Sparse (6 frames), early-half (first 50% of video), and late-half (second 50% of video). For each query, we extract the token-level probability distribution over answer options and compute three confidence metrics: maximum probability ($p_{\max}$), margin, and normalized entropy.
@@ 5.1.2 Calibration Analysis
Figure 3: Takeaway: In-distribution, confidence-based abstention works. Baseline performance on 18 frames ($n=300$). (a) Risk-Coverage: sweeping $\varepsilon$ produces a smooth, monotone curve—higher thresholds reliably reduce error. Starred point: $\varepsilon=0.71$ achieves 9.4% risk at 63.7% coverage. (b) Reliability diagram (all predictions): bars show accuracy per confidence bin. The model is well-calibrated—predicted confidence matches actual accuracy. Most predictions cluster at high confidence (0.9–1.0 bin has $n=189$). (c) Expected Calibration Error (ECE) vs. threshold: calibration error drops from 0.067 to 0.018 as we become more selective.
@@ 5.2 Evidence Degradation
Figure 4: Overconfidence under evidence degradation. Question: “How does the brown dog keep the white dog down?” Options: A) change hand, B) hold the dog with its paws, C) walks around, D) using leash, E) lie down on chair. Answer: B. With 18 frames, the model answers correctly (B) with confidence 1.00. With only 6 frames (Sparse condition), the model answers incorrectly (A) yet still reports confidence 0.70. The question requires observing sustained behavior across time; sparse sampling misses critical moments, but the model fails to recognize that its evidence is insufficient.
@@ 5.2.1 The Overconfidence Problem
Figure 5 captures the core finding directly: a 67% reduction in evidence produces less than 6% change in confidence. Self-reported confidence drops only 3.3%, and logprob-derived $p_{\max}$ actually increases by 0.6%. The gap between evidence reduction and confidence reduction is the problem. A well-calibrated epistemic system would show commensurate contraction; instead, confidence remains high despite reduced observability.
Figure 5: Per-item confidence comparison under evidence degradation (18 $\to$ 6 frames) on matched question instances. Each point is one question; both panels show the same $n=292$ items. Panel (A) appears sparser because self-reported confidence is discretized (0.6, 0.7, 0.9, 1.0, etc.), causing points to stack—darker regions indicate more overlapping points. If confidence tracked evidence, points would fall toward the dashed line (proportional 67% reduction). Instead, points cluster near the diagonal, showing confidence is insensitive to evidence quality.
[…]
Figure 6: Takeaway: Even with 67% less evidence, high-confidence mass barely decreases—and $p_{\max}$ actually increases—on the same questions. Left/middle: CDFs with a fixed high-confidence threshold ($c\geq 0.9$). Right: paired-dots summary of high-confidence mass. Self-reported and logprob confidence are computed on matched question instances (same question_id, same evidence condition) via separate model calls; statistics use the matched intersection where both outputs are valid ($n=295$ at 18 frames, $n=292$ at 6 frames).
@@
Figure 7: Risk-Coverage comparison under evidence degradation. Coverage is the fraction of predictions accepted (not abstaining); Risk is the error rate among accepted predictions. The Sparse curve (6 frames, dashed) shifts leftward and upward from the original (18 frames, solid). At any fixed $\varepsilon$, the 6-frame regime has lower coverage but similar conditional risk. Compressed (dotted) has minimal impact.
@@ 5.3 Logprob Confidence
Sparse (6 frames) 81.3% 0.876 0.974 0.771 0.282
Early-half (6 frames) 77.3% 0.861 0.961 0.744 0.319
Late-half (6 frames) 78.7% 0.872 0.979 0.772 0.302
@@ 5.3.1 Comparison with Self-Reported Confidence
Table 5 directly compares self-reported confidence (from JSON output) with logprob-derived $p_{\max}$ on matched question instances (same question_id, same evidence condition) via separate model calls. The comparison reveals that logprob confidence is systematically higher and less responsive to evidence degradation. Figure 8 visualizes the risk-coverage curves for both confidence methods under evidence degradation.
Figure 8: Risk-coverage curves comparing self-reported confidence (left) and logprob-derived $p_{\max}$ (right) under evidence degradation. Computed on matched question instances via separate model calls. Both methods exhibit leftward shift when frames are reduced from 18 to 6. However, the gap between conditions is smaller for logprob confidence, indicating that the model’s token-level decision distribution is even less sensitive to evidence quality than self-reported confidence.
@@
Mean 0.832 0.804 0.870 0.876
Median 0.900 0.900 0.969 0.975
$\Delta$ (degradation) -3.3% ++0.6%
Table 5: Self-reported vs. logprob-derived confidence under evidence degradation. Computed on matched question instances (same question_id, same evidence condition) via separate model calls; statistics use the matched intersection where both outputs are valid ($n=295$ at 18 frames, $n=292$ at 6 frames). Self-reported confidence drops 3.3% (0.832 $\to$ 0.804) when frames are reduced. Logprob $p_{\max}$ actually increases slightly (0.870 $\to$ 0.876). The model’s token-level decision distribution is even less sensitive to evidence quality than its self-reported confidence.
@@ 5.3.2 Interpretation
  • Self-reported confidence drops 3.3% under degradation (0.832 $\to$ 0.804)
  • Logprob $p_{\max}$ increases 0.6% (0.870 $\to$ 0.876)
@@
The temporal ablation conditions (early-half and late-half) reinforce this conclusion. For questions whose answers depend on events in the missing temporal segment, these conditions provide semantically insufficient evidence by construction. Yet $p_{\max}$ remains above 0.86 and median $p_{\max}$ exceeds 0.96 in both cases. Because temporal ablation guarantees removal of one half of the clip, stability of $p_{\max}$ under early-half and late-half conditions suggests insensitivity is not explained solely by redundant sampling—the model genuinely does not recognize when critical temporal context is absent.
@@ 6.1 Summary of Findings
Third, the logprob analysis confirms that overconfidence is fundamental to the model’s representations, not an artifact of the self-reporting interface. Logprob-derived $p_{\max}$ actually increases slightly under evidence degradation (0.870 $\to$ 0.876), while self-reported confidence at least decreases modestly (0.832 $\to$ 0.804). The model’s token-level decision distribution over answer options does not become more diffuse when evidence quality degrades.
@@ 6.3.1 What These Experiments Support
These three facts imply a warrant-based formulation in which reported confidence $p$ should satisfy $p\leq\zeta(e)+\epsilon$ for some evidence-derived bound $\zeta(e)$. The Sparse condition failure is precisely the violation this contract would rule out. Confidence remains high when the evidence channel is weaker.
@@ 6.3.6 Observability Sensitivity Diagnostic
$\Pr(\text{conf}\geq 0.9)$ 64.1% (189/295) 54.8% (160/292)
$\Pr(\text{wrong}\mid\text{conf}\geq 0.9)$ 9.0% 8.8%
Mean confidence 0.832 0.804
@@
Confidence does not contract commensurate with evidence loss. Despite reducing frames from 18 to 6, median confidence remains identical at 0.900. The high-confidence rate drops modestly (64.1% to 54.8%), but the error rate among high-confidence predictions is nearly identical (9.0% vs 8.8%). Confidence is not evidence-conditioned. The model maintains high confidence despite reduced observability, and the modest selectivity increase (coverage drops 14% at fixed $\varepsilon$) is not proportional to the 67% evidence reduction.
@@ Appendix B Per-Category Degradation
We analyze degradation patterns across question categories (Causal: CW+CH, Temporal: TN+TC+TP, Descriptive: DO+DL+DC) at three operating points to prevent cherry-picking. Note that the overall coverage reported in Table 3 (e.g., 53.7% at $\varepsilon=0.71$ under Sparse) is the average across these three 100-item strata; the per-category breakdown below explains the aggregate behavior.
@@ Appendix C Temporal Ablation
For each video, we generate two 6-frame evidence packets: early-half (frames sampled uniformly from the first 50% of the clip) and late-half (frames sampled uniformly from the second 50%). This design targets temporal and causal questions, which often require observing sequences of events spanning the full video. By restricting frames to one half, we remove evidence about events occurring in the other half—a more controlled intervention on semantic content than uniform subsampling, which may still capture key moments regardless of density.
@@
Sparse (6 uniform) 97.3% 27.4% 72.6% 0.786
Early-half (0–50%) 97.3% 25.7% 74.3% 0.792
Late-half (50–100%) 97.7% 25.3% 74.7% 0.800
@@
Sparse (6 uniform) 53.7% 9.3% 90.7% 0.926
Early-half (0–50%) 52.3% 8.9% 91.1% 0.923
Late-half (50–100%) 51.7% 10.3% 89.7% 0.929
@@
Table 10: Temporal Ablation results. At $\varepsilon=0$, early-half and late-half conditions perform similarly to uniform 6-frame sampling (25–26% risk vs 27%). At $\varepsilon=0.71$, all 6-frame conditions converge to similar coverage (51–54%) with mean confidence remaining above 0.92. The model does not differentiate between early and late evidence despite the procedural removal of temporal context.
@@
Temporal Ablation conditions behave nearly identically to uniform 6-frame subsampling. At $\varepsilon=0$, early-half (25.7% risk) and late-half (25.3% risk) perform comparably to Sparse (27.4% risk). At $\varepsilon=0.71$, all three 6-frame conditions converge to similar coverage (51–54%) and mean confidence ($\sim$0.92). The model does not “notice” whether it is seeing the first or second half of the video. Confidence remains high regardless of which temporal segment is provided.
@@
This finding strengthens our claim that confidence does not track evidence completeness. Unlike uniform subsampling, which may preserve key moments by chance, temporal ablation guarantees removal of one half of the clip’s temporal context. For questions whose answers depend on late events (e.g., “what does X do after Y?”), early-half evidence is semantically insufficient. Yet the model maintains similarly high confidence in both conditions. This suggests overconfidence is not merely an artifact of temporal redundancy in the dataset, but reflects a fundamental insensitivity to evidence availability.
Appendix D Formal Metric Definitions
This appendix provides complete mathematical definitions for all evaluation metrics used in this work.
D.1 Selective Prediction Metrics
Let $P$ denote the full set of $n$ predictions, and let $A_{\varepsilon}\subseteq P$ be the subset of predictions accepted at confidence threshold $\varepsilon$—i.e., predictions where the model’s confidence $c_{i}\geq\varepsilon$.
Coverage.
The fraction of inputs for which the model provides an answer:
$\text{Coverage}(\varepsilon)=\frac{|A_{\varepsilon}|}{|P|}$ (4)
Risk.
The error rate among accepted predictions:
$\text{Risk}(\varepsilon)=\frac{1}{|A_{\varepsilon}|}\sum_{i\in A_{\varepsilon}}\mathbf{1}[\hat{y}_{i}\neq y_{i}]$ (5)
where $\hat{y}_{i}$ is the predicted answer and $y_{i}$ is the ground truth.
Conditional Accuracy.
The complement of risk:
$\text{Acc}_{\text{cond}}(\varepsilon)=1-\text{Risk}(\varepsilon)=\frac{1}{|A_{\varepsilon}|}\sum_{i\in A_{\varepsilon}}\mathbf{1}[\hat{y}_{i}=y_{i}]$ (6)
D.2 Calibration Metrics
Expected Calibration Error (ECE).
Predictions are partitioned into $B$ equal-width bins by confidence. Let $B_{b}$ denote the set of predictions in bin $b$. ECE measures the weighted average gap between confidence and accuracy:
$\text{ECE}=\sum_{b=1}^{B}\frac{|B_{b}|}{n}\left|\text{acc}(B_{b})-\text{conf}(B_{b})\right|$ (7)
where:
$\text{acc}(B_{b})$ $=\frac{1}{|B_{b}|}\sum_{i\in B_{b}}\mathbf{1}[\hat{y}_{i}=y_{i}]$ (8)
$\text{conf}(B_{b})$ $=\frac{1}{|B_{b}|}\sum_{i\in B_{b}}c_{i}$ (9)
We use $B=10$ bins throughout. A perfectly calibrated model has $\text{ECE}=0$.
D.3 Logprob-Derived Confidence Metrics
Given raw log probabilities $\{\ell_{A},\ell_{B},\ell_{C},\ell_{D},\ell_{E}\}$ from the model’s token distribution over the five answer options, we first normalize via softmax:
$p_{i}=\frac{\exp(\ell_{i})}{\sum_{j\in\{A,B,C,D,E\}}\exp(\ell_{j})}$ (10)
From this distribution $\mathbf{p}=(p_{A},p_{B},p_{C},p_{D},p_{E})$, we derive three confidence metrics:
Maximum Probability ($p_{\max}$).
The probability assigned to the most likely answer:
$p_{\max}=\max_{i\in\{A,B,C,D,E\}}p_{i}$ (11)
Higher $p_{\max}$ indicates the model concentrates probability mass on a single option.
Margin.
The gap between the top two probabilities:
$\text{Margin}=p_{\max}-p_{\text{second}}$ (12)
where $p_{\text{second}}$ is the second-highest probability. Larger margins indicate more decisive predictions.
Normalized Entropy.
Uncertainty measured as the ratio of actual entropy to maximum entropy:
$\text{Entropy}_{\text{norm}}=\frac{H(\mathbf{p})}{H_{\max}}=\frac{-\sum_{i}p_{i}\log p_{i}}{\log 5}$ (13)
where $H_{\max}=\log 5$ is the entropy of a uniform distribution over 5 options. Values near 0 indicate high confidence (concentrated distribution); values near 1 indicate high uncertainty (uniform distribution).
Appendix E Prompt Templates
We use two different prompts for self-reported confidence (JSON output) and logprob-derived confidence (letter-only output).
E.1 Self-Reported Confidence Prompt